OpenStudy (anonymous):

find the equation of the line tangent to the graph of y=3x-cosx at x=0.

OpenStudy (rogue):

\[y = 3x - \cos x, y(0) = -1\]\[y' = 3 + \sin x, y'(0) = 3 + sin (0) = 3\]\[Y_{tangent} = 3x - 1\]